# Quantum Mikheyev–Smirnov–Wolfenstein effect

### v1

@@ Appendix A Massive Weyl neutrino propagators in matter
Note that $\eta_{a}(\mathbf{x},t)$ in Eq. (A.3) and analogous $\eta_{a}^{*}(\mathbf{x},t)$ obey the canonical equal time anticommutation relation $\left\{\eta_{a}(\mathbf{x},t),\eta_{a}^{*}(\mathbf{y},t)\right\}=\delta(\mathbf{x}-\mathbf{y})$ provided that $\left\{a_{\pm}(\mathbf{p}),a_{\pm}^{\dagger}(\mathbf{q})\right\}=\delta(\mathbf{p}-\mathbf{q})$ and all other anticommutators vanish. Thus, we have constructed the canonical quantization of Weyl neutrinos in background matter.
@@
where $\theta(t)$ is the Heaviside step function. Using Eqs. (A.3), (A.4), and (A.6), as well as the identity
@@
-- $\lambda_{a+}^{2}\left(\frac{1}{p_{0}+E_{a+}-\mathrm{i}0}+\frac{A_{a+}^{2}}{p_{0}-E_{a+}+\mathrm{i}0}\right)\bigg{]}\bigg{\}},$ (A.10)
@@
In this Appendix, we calculate the Fourier transform of the propagator of a single massive Weyl neutrino in matter. We use the corresponding propagators in Appendix A.
@@
$I_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0},\quad J_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0},$ (B.2)
@@
$I$ =${}_{1}=\frac{1}{(2\pi)^{2}}\int_{0}^{+\infty}\rho\mathrm{d}\rho\int_{-\infty}^{+\infty}\mathrm{d}z\frac{(E_{a-}+q+g_{a})e^{\mathrm{i}zL}}{2E_{a-}(E-\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}+\mathrm{i}0)}.$ (B.3)
@@
Considering the limit of the great distance of the neutrino propagation, one has that $J_{1}\approx(\bm{\sigma}\hat{L})I_{1}$.
@@
In Eq. (B.11), we show the limit of great $L$. One can see in Eqs. (B.6) and (B), that $|I_{1,3}|\gg|I_{2,4}|$ for ultrarelativistic particles with $E\gg m_{a}$.
According to Eq. (2.10), the matrix element is proportional to the Fourier transform of the propagator in Eq. (B.1) averaged over the lepton wavefunctions. For simplicity, we neglect the mixing matrix in Eq. (2.10). Considering ultrarelativistic left leptons in initial and final states with $\mathbf{p}\parallel\mathbf{L}$, as well as using Eqs. (B.6), (B.9), (B), and (B.11), one has that

### v2

@@ Appendix A Massive Weyl neutrino propagators in matter
Note that $\eta_{a}(\mathbf{x},t)$ in Eq. (A) and analogous $\eta_{a}^{*}(\mathbf{x},t)$ obey the canonical equal time anticommutation relation $\left\{\eta_{a}(\mathbf{x},t),\eta_{a}^{*}(\mathbf{y},t)\right\}=\delta(\mathbf{x}-\mathbf{y})$ provided that $\left\{a_{\pm}(\mathbf{p}),a_{\pm}^{\dagger}(\mathbf{q})\right\}=\delta(\mathbf{p}-\mathbf{q})$ and all other anticommutators vanish. Thus, we have constructed the canonical quantization of Weyl neutrinos in background matter.
@@
where $\theta(t)$ is the Heaviside step function. Using Eqs. (A), (A.4), and (A.6), as well as the identity
@@
$-\lambda_{a+}^{2}\left(\frac{1}{p_{0}+E_{a+}-\mathrm{i}0}+\frac{A_{a+}^{2}}{p_{0}-E_{a+}+\mathrm{i}0}\right)\bigg{]}\bigg{\}},$ (A.10)
@@
In this Appendix, we calculate the Fourier transform of the propagator of a single massive Weyl neutrino in matter. The corresponding propagators are taken from Appendix A.
@@
$I_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0},\quad J_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0}.$ (B.2)
@@
$I_{1}=\frac{1}{(2\pi)^{2}}\int_{0}^{+\infty}\rho\mathrm{d}\rho\int_{-\infty}^{+\infty}\mathrm{d}z\frac{(E_{a-}+q+g_{a})e^{\mathrm{i}zL}}{2E_{a-}(E-\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}+\mathrm{i}0)}.$ (B.3)
@@
Considering the limit of large neutrino propagation distance, one has that $J_{1}\approx(\bm{\sigma}\hat{L})I_{1}$.
@@
In Eq. (B), we show the limit of great $L$. One can see in Eqs. (B.6) and (B), that $|I_{1,3}|\gg|I_{2,4}|$ for ultrarelativistic particles with $E\gg m_{a}$.
According to Eq. (2.10), the matrix element is proportional to the Fourier transform of the propagator in Eq. (B.1) averaged over the lepton wavefunctions. For simplicity, we neglect the mixing matrix in Eq. (2.10). Considering ultrarelativistic left leptons in initial and final states with $\mathbf{p}\parallel\mathbf{L}$, as well as using Eqs. (B.6), (B), (B), and (B), one has that